OpenStudy (anonymous):

The gravitational attraction between two masses varies inversely as the square of the distance between them. If the attraction force between them is 75 lbs. when the bodies are eight feet apart, find the attraction when the masses are twelve feet apart. inverse variation help?

OpenStudy (anonymous):

ha I had to look at hte name bc of all hte pic copying lol

sam (.sam.):

\[\huge F=\frac{GM_{1}M_{2}}{r^{2}}\]

sam (.sam.):

assuming masses are equal, and G is constant, \[75=\frac{M}{8^2}\] M=4800 ---------------------------- \[F=\frac{4800}{12^2}=33.33\]

OpenStudy (agreene):

@.sam if you're going to use Newton's equation (which was proven wrong in the 1930's) at least know that G is a constant. \[G\approx 6.674 \times 10^{-11}N(m/kg)^2\] In this case, actually using that equation is erroneous, due to the fact the question provides you with the equation they want you to use (which u did end up using because you were unaware of G)

sam (.sam.):

@becca94 is this correct? @agreene the G will cancel later

OpenStudy (anonymous):

that's a lot...I don't know... :(

OpenStudy (anonymous):

the answer was 33.3 :)

sam (.sam.):

Yeah look at my work